Subject: APO 3.5-m users committee minutes, May 11, 2009
From: Suzanne Hawley
Submitted: Fri, 15 May 2009 12:16:25 -0700 (PDT)
Message number: 1132
(previous: 1131,
next: 1133
up: Index)
APO 3.5-m Users Committee Phonecon, 5/11/09
Attending: Suzanne Hawley, Scott Anderson, Jon Fulbright, Mark
Klaene, Bill Ketzeback, Bruce Gillespie, Michael Strauss, Remy
Indebetouw, and Guy Stringfellow (for John Bally)
Absent: Al Harper, John Bally, Jon Holtzman
**********************************
User feedback, comments from institutional representatives:
Colorado (Guy Stringfellow) - Guy mentioned that he had heard of user
complaints that TripleSpec beam switching on bright targets isn't fast
enough. Bill Ketzeback said that this might be due to the instrument
rotator not being accurately zeroed, and/or problems with TripleSpec
instrument balance. We are planning to perform on-sky testing to try
to characterize this during upcoming engineering time.
Princeton (Michael Strauss) - Michael said things were fine, but
mentioned that PU users had also been impacted by recent incidences of
problems with TripleSpec beam switching.
Johns Hopkins (Jon Fulbright) - Jon had nothing new to report.
Washington (Scott Anderson) - Scott had nothing new to report.
Chicago (Al Harper) - no report.
New Mexico State (Jon Holtzman) - no report.
Virginia (Remy Indebetouw) - Remy reported that there was "general
happiness" among UVa users.
**********************************
Discussion of telescope/instruments report:
Mark said that we've seen some wind, dust, clear skies, and high fire
danger in the past month. Dave Woods is retiring as the 3.5-m's
electronics technician next year, and his replacement, Ed Leon, is
starting this week. The telescope and instruments have been generally
running well. The APOLLO laser is operational again. The direct-
drive replacement project is on track for the late-summer shutdown.
Bruce mentioned that if any ARC institution is planning to bring large
groups of students to the site this summer, there may be traffic
congestion due to the large amount of engineering work planned for the
3.5-m and SDSS telescopes between June and mid-September--it would be
best if student groups came to APO outside of this time frame.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 4/11/09
through 5/6/09
Mark Klaene
0) Overview
Weather continues to be good with moderating winds but atypical spring
cirrus clouds. The lower part of the exterior of the dome and arcade
have been repainted. Annual forest fire prevention activities are
underway, e.g., a fire break around the site has been recut and four
truckloads of forest detritus have been hauled off.
An Electronics Technician was hired and will start this Wednesday, who
will be the replacement for Dave Woods when he retires early next
year. We have so far not received any promising applications for the
posted Telescope Engineer position.
1) Telescope
Telescope operations have been smooth, with some excellent seeing
reported and generally good telescope pointing and tracking
performance. No significant telescope problems arose this past month.
2) Instruments
DIS, SPIcam, and the Echelle are fully operational, and NIC-FPS has
been operational with no problems. The new NIC-FPS fan-out board
needs to be populated, tested at UVa and CU, and then a plan for its
installation will be formulated--CU commitments to HST/COS will delay
the fan-out board installation until after the summer.
TripleSpec has been operating with a borrowed Leach power supply so we
can monitor the array power dropouts. The frequency of these dropouts
has been significantly reduced. Some temperature read-out problems
are still occurring, however.
Agile is still available at NA2. Its move to the TR2 port has been
delayed due to scheduling conflicts but we expect that it will be
moved during the June engineering time. There is a new TUI release for
AGILE control. There has also been more study and data collected
regarding Agile bias shifts and overscan areas; the Agile documentation
is being updated and should be consulted prior to using the instrument.
The APOLLO laser is currently offline due to a problem which may be
remedied this weekend. A precision gravimeter has been installed in
the cone room beneath the telescope, and it is producing useful
telescope positional information. The aircraft transponder receiver
array is also in operation. The APOLLO team is making plans for future
work with the Lunar Reconnaissance Orbiter.
3) CIF projects
The direct-drive project is well on its way. Machining schedules are
very tight, but we are still on schedule for a late-August or early-
September installation.
**********************************
On-site training:
Suzanne said that we have always encouraged student visits to the
telescope, particularly as part of their orientation to the facilities
and for training on how to use them. It was recently asked whether
the 3-night training requirement applies to postdocs who are new
users. The training policy
(see http://www.apo.nmsu.edu/arc35m/TrainingInfo_ARC35m.html)
stipulates that all new 3.5-m telescope users are required to spend
3 nights at the telescope for orientation and training. Students need
to be accompanied by their faculty advisor or a designated experienced
observer; new Ph.D.-level observers can receive their training/
orientation by themselves. If an observer, one who hasn't had the on-
site orientation, plans to run the telescope remotely, he or she will
need to be assisted at all times by an astronomer who has been trained
on-site. This policy has been in place since the telescope was first
opened, and is intended, in part, to promote the hands-on experience
of observing for new telescope users. Every few years, questions
about this policy have arisen, and in this instance, the Users
Committee members endorsed the continuation of the training policy.
**********************************
Instrument updates:
Suzanne gave an update on the status of the new instrument
initiatives. Eric Burgh is planning to issue a report with
recommendations for upgrading the echelle spectrograph sometime in
June. Cynthia Froning is also preparing a proposal for a new visible-
imaging capability for the telescope; her report is expected in late
June. Jon Holtzman has been talking with WIYN about the possibility
of a loan of the QUOTA imager--Suzanne will send the committee a link
that describes QUOTA. Also at various stages of discussion are three
other possible new instrumentation projects: a wide-field NIR imager
from Wyoming, a high-res, fiber-fed, NIR spectrograph from UFla, and
upgrades to the Goddard Fabry-Perot instrument.
**********************************
Summer shutdown:
Mark announced that we are having two shutdowns this year, one short
(1-week) in July to wash M1 and service its mirror cell, and a longer
one (4-week) in mid-August through mid-Sept to install the new direct-
drive actuators on altitude and azimuth, with their new axis
controllers. Suzanne added that the latter shutdown is ambitious, and
risks to the schedule are possible that could affect post-shutdown
observing schedules. If the installation of the new drives encounters
unexpected difficulties, fall-back plans are in place.
**********************************
ACTION ITEMS:
[open from previous months]:
none
[new actions from this meeting]:
none
**********************************
Next meeting: The next Users Committee phonecon will be on Monday, 8
June, at 8:30 AM Pacific Time. The agenda and other materials will be
sent to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1132 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO